### v1

# Probing supersolidity through excitations in a spin-orbit-coupled Bose-Einstein condensate

Abstract
Spin-orbit-coupled Bose-Einstein condensates are a flexible experimental platform to engineer synthetic quantum many-body systems. In particular, they host the so-called stripe phase, an instance of a supersolid state of matter. The peculiar excitation spectrum of the stripe phase, a definite footprint of its supersolidity, has so far remained out of experimental reach. We achieve in situ imaging of the stripes and directly observe both superfluid and crystal excitations. We investigate superfluid hydrodynamics and reveal a stripe compression mode, thus demonstrating that the system possesses a compressible crystalline structure. Through the frequency softening of this mode, we locate the supersolid transition point. Our results establish spin-orbit-coupled supersolids as a platform of choice to investigate supersolidity and its rich dynamics.
Supersolidity is an exotic phase of matter characterized by two spontaneously-broken continuous symmetries, translational symmetry and U(1) gauge symmetry [1, 2]. They are associated to two independent order parameters and manifest themselves as a modulated density, reminiscent of a solid, and a global phase coherence, characteristic of a superfluid [3, 4, 5]. This dual nature leads to a remarkably rich excitation spectrum with both superfluid and crystal excitations [6], including two gapless Goldstone modes. After the first realizations of supersolids in quantum gases [7, 8, 9, 10, 11], exploring their dynamics has thus been the focus of experimental research in recent years [12, 13, 14, 15, 16, 17, 18, 19] as it provides an unambiguous signature of supersolidity. Bose-Einstein condensates (BECs) with spin-orbit coupling, where the internal states of the particles are linked to their momentum through optical coupling [20], have been shown to host a supersolid state of matter in the form of the so-called stripe phase. Due to its conceptual simplicity, it can be described analytically and is thus ideally suited to study the exotic properties of supersolids [21, 22, 20, 23, 24]. However, only very fragile realizations of the stripe phase have been demonstrated so far, and the contrast of the stripes, i.e., of the density modulation, was vanishingly small [8, 25]. As a result, the excitations of this phase have been widely believed to be experimentally inaccessible and a debate has arisen over their nature [26]. It was fueled by the observation that other platforms fall into two distinct categories: supersolids that support phonon modes [13, 15], and supersolids that are stiff [12], meaning that the periodicity of their density modulation is fixed. This led to the following question: are spin-orbit-coupled supersolid stripes stiff or do they host phonon modes?
In this work, we experimentally explore the excitation spectrum of the stripe phase and show that spin-orbit-coupled supersolids have a compressible crystal structure and are not stiff. By harnessing the tunable interaction properties of ultracold potassium, we overcome the limitations of previous approaches and realize a stable stripe phase that persists at coupling regimes with strong modulation contrast. While the modulated density pattern could only be probed indirectly in previous experiments [8, 25], here we achieve in situ imaging of the stripes using a matter-wave magnification scheme. By exciting the collective modes of the system, we are able to address both superfluid and crystal-like aspects of supersolids. We show the existence of particle flow between the stripes and observe a stripe compression mode, i.e., an oscillation of the modulation period, which unambiguously proves the compressibility of the crystalline structure. Furthermore, we locate the phase transition from the supersolid to an unmodulated superfluid through frequency softening of the compression mode.
I.   Supersolid spin-orbit-coupled condensates
[…]
In spin-orbit-coupled BECs, supersolidity stems from the unconventional dispersion relation and the momentum-dependent interactions that arise due to the mixing of spin and spatial degrees of freedom [21, 20, 22, 23]. The dispersion relation displays two minima in momentum space [28], which results in two possible phases at the many-body level, sketched in Fig. 1(a). The plane-wave phase, which appears when a single minimum is occupied, has a homogeneous density profile. In contrast, the stripe phase, which occurs when both minima are occupied, has a spontaneously modulated density profile while it retains the spatial phase coherence of the condensate [21, 22, 20, 23], and constitutes a supersolid [6]. The experimental challenge is to create a situation where the stripe phase occupies a significant portion of the phase diagram [20]. To quantitatively investigate the stability of the stripe phase and find a suitable regime to probe supersolid excitations, we have developed an analytical model.
This model considers the dressed states at the two minima of the dispersion relation, which emerges when two different states $\downarrow$ and $\uparrow$ (e.g., internal atomic states [20, 25], or different bands of an optical lattice [8]) are coupled through an optical field that transfers momentum to the atoms. A pair of laser beams in a two-photon Raman configuration with frequency difference $\Delta\omega_{\mathrm{R}}$, two-photon Rabi frequency $\Omega$ and two-photon detuning $\delta$ is used, as depicted in Fig. 1(b). It leads to a momentum transfer $2\hbar k_{\mathrm{R}}$ along the $x$ direction $\mathbf{e_{x}}$, with $\hbar k_{\mathrm{R}}$ the recoil momentum unit and $\hbar=h/(2\pi)$ the reduced Planck’s constant. The atom-photon dressed states have a momentum-dependent spin composition, which is set by the polarization parameter $P=\tilde{\delta}/\tilde{\Omega}$, where $\tilde{\delta}=\delta-2\hbar k_{\mathrm{R}}k/m$ and $\tilde{\Omega}=\sqrt{\Omega^{2}+\tilde{\delta}^{2}}$ are the generalized detuning and Rabi frequency. Here $m$ is the atomic mass and $\hbar k$ is the momentum along $x$ in the frame rotating at $\Delta\omega_{\mathrm{R}}$ and co-moving with the Raman coupling fields [29]. In this frame, the single-particle Hamiltonian of the system takes the translationally invariant form
$\mathcal{H}_{\mathrm{kin}}=\frac{\hbar^{2}}{2m}(\mathbf{k}+k_{\mathrm{R}}\sigma_{z}\mathbf{e_{x}})^{2}-\frac{\hbar\delta}{2}\sigma_{z}+\frac{\hbar\Omega}{2}\sigma_{x},$ (1)
where $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices. For $\hbar\Omega<4\,E_{\mathrm{R}}$, where $E_{\mathrm{R}}=\hbar^{2}k_{\mathrm{R}}^{2}/2m$ is the Raman recoil energy, the dispersion relation of the lower dressed state features two minima, see Fig. 1(c). The system behaves as a mixture of condensates in the left ($\ell$) and right ($r$) wells that are non-orthogonal. Thus, a modulated density profile emerges due to interference between them when both $\ell$ and $r$ are simultaneously occupied and miscible [28, 20]. Their overlap increases with the spin-orbit coupling strength, leading to a contrast of the density modulation that, away from the transition, scales linearly with $\Omega$ [29].
Figure 2: In situ observation of supersolid stripes. (a) Single shot experimental image of the in situ density profile of the cloud at $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$ measured with matter-wave optics. The image is rescaled to represent the expected cloud aspect ratio before matter-wave magnification along $x$. Top right: corresponding density profile integrated along the stripes. The black line denotes a modulated Gaussian fit with stripe spacing $d$. Bottom right: histogram showing the spatial phase of the modulation $\phi$ over 350 realizations. (b) Spacing $d$ of the stripes vs. Raman coupling $\Omega$ measured from in situ (blue circles) and time-of-flight (gray squares, see [29]). Solid line: single-particle theory prediction, obtained from the inverse of the momentum difference of the minima of the dispersion relation (shown on the right for different spin-orbit coupling strengths $\Omega$). Errorbars correspond to one standard error of the mean (eom). The blue shaded area shows the systematic error of the in situ data from the magnification calibration.
Our effective field theory is based on the effective interactions of the optically-dressed states, which are given by the polarization at the dispersion minima [29]. We name it the mixture model because it builds on the mixture picture introduced by [28, 20]. It delivers analytic expressions for both ground state properties and collective excitations [29] and highlights the conceptual simplicity of spin-orbit-coupled supersolids. Figure 1(d) shows the effective interaction parameters according to our mixture model for a gas with bare scattering lengths $a_{\uparrow\uparrow}$, $a_{\downarrow\downarrow}$, and $a_{\uparrow\downarrow}$ at detuning $\delta=0$. Increasing $\Omega$ reduces the effective intra-well interactions $a_{\ell\ell}$, $a_{rr}$ and increases the inter-well ones $a_{\ell r}$. As a result, the miscibility of the system is reduced and it becomes energetically favorable to occupy a single minimum of the dispersion relation as soon as $a_{\ell r}>a_{\ell\ell}$ or $a_{\ell r}>a_{rr}$. The crystal-like modulation is then lost and the system transitions to the plane-wave state. This relation between interactions and coupling strength demonstrates that obtaining a stripe phase at large couplings requires $a_{\uparrow\downarrow}\ll a_{\uparrow\uparrow},a_{\downarrow\downarrow}$. This situation is not available in atomic species like the 87Rb used in previous works, where the small scattering length difference yields a very fragile supersolid phase with a contrast of the density modulation $C<5\%$ [20, 30, 31, 25]. However, we can achieve it by exploiting two-component potassium BECs, where the interactions are tunable using Feshbach resonances [32]. The complete phase diagram of the system with 41K at 51.7 G in the parameter space of $\Omega$ and $\delta$ is depicted in Fig. 1(e). The stripe phase appears as a dome that is not centered at $\delta=0$ due to the setting $a_{\downarrow\downarrow}<a_{\uparrow\uparrow}$. The phase diagram shows the contrast $C$ according to the variational model of Refs. [23, 27], in excellent agreement with the black dashed lines, which are analytical predictions from our mixture model. This constitutes an unusually large, stable, and high-contrast supersolid stripe phase and provides an excellent regime to perform our experiments.
II.   In situ observation of supersolid stripes
We realize the supersolid stripe phase with a Bose-Einstein condensate of 41K atoms in states ||$\ket{\downarrow}\equiv\ket{F=1,m_{F}=0}$ and ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$. By setting the magnetic field to $B=51.7\pm 0.1$ G, between an intra-spin and an inter-spin Feshbach resonance, we tune the bare state interactions to $(a_{\uparrow\uparrow},a_{\downarrow\downarrow},a_{\uparrow\downarrow})=(115\pm 10,65\pm 1,40\pm 8)\,a_{0}$, where $a_{0}$ is the Bohr radius [32]. We couple the two states with two Raman laser beams of orthogonal polarizations at the tune-out wavelength $\lambda_{\mathrm{R}}=768.97$ nm, for which scalar potentials cancel. We work at a zero two-photon Raman detuning $\hbar\delta=0.00\pm 0.02\,E_{\mathrm{R}}$ (gray band in Fig. 1(e)), where the phase transition point is independent of the atomic density for our parameter regime [33]. The Raman beams have a relative angle of $\theta=94\pm 1^{\circ}$, which leads to a momentum transfer along the $x$ direction of $2\hbar k_{\mathrm{R}}=4\pi\hbar\mathrm{sin}(\theta/2)/\lambda_{\mathrm{R}}$ [29].
[…]
To investigate the periodicity of the stripes, we extract the stripe spacing from in situ images taken at different spin-orbit coupling strengths $\Omega$. Figure 2(b) shows that the stripe spacing $d$ increases significantly with $\Omega$, which can be understood from a momentum-space picture. The stripe spacing is given by the momentum difference $\Delta k$ of the two dressed states through $d=2\pi/\Delta k$ which, at equilibrium, corresponds to the difference between the two minima of the dispersion relation (see right panel). It can also be inferred from spin-resolved momentum-space images obtained through time-of-flight expansion. Contrary to the in situ one, this analysis is model dependent [29], resulting in a spacing $\pi/k_{\mathrm{R}}$ at the zero-coupling point by construction. The data is shown as gray squares in Fig. 2(b). Within their respective uncertainties, both in situ and time-of-flight data agree with the theory prediction. Thus, our observations demonstrate that the period of spin-orbit-coupled supersolids is not externally imposed by the wavelength of the coupling laser. This variable crystalline pattern stands in contrast to cavity-mediated supersolidity, where the modulation period is fixed by the light wavelength [12].
III.   Probing superfluid hydrodynamics through excitations
Figure 3: Crystal Goldstone mode and superfluid hydrodynamics. (a) Population imbalance between side stripes $\eta$ vs. displacement of the central stripe $D$ with respect to the center of mass of the cloud (see inset and [29]). For a harmonically-trapped system with superfluid flow, where a sliding of the density modulation is compensated by particle exchange between the stripes, both quantities are linearly correlated and provide evidence for the crystal Goldstone mode. (b) Top panel: Breathing mode of the overall cloud, observed through its size $\sigma$ vs. time. Solid line: sinusoidal fit used to extract the breathing frequency $\omega_{\mathrm{B}}$, yielding $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1\sim\sqrt{5/2}$ with $\omega_{\mathrm{D}}$ the dipole mode frequency, as expected from superfluid hydrodynamics. Bottom panel: number of stripes in the supersolid cloud. It dynamically changes during half of a breathing oscillation (see shaded area in top panel), supporting the existence of superfluid flow across the system. Errorbars denote one standard eom.
Supersolids are characterized by a superfluid flow through the crystalline structure that maintains the global phase coherence of the system [5, 1]. While previous experiments demonstrated phase coherence in spin-orbit-coupled supersolids using Talbot interferometry [25], our in situ images allow us to investigate instead the superfluid hydrodynamic properties of the system. To this end, we followed the procedure developed for supersolids in dipolar gases [14] and studied the population imbalance between stripes. Figure 3(a) shows the imbalance between side stripes $\eta$ as a function of the displacement $D$ of the central stripe with respect to the center of mass of the cloud (see illustration of $D$ and $\eta$ in the inset and [29]). We see a clear linear correlation between displacement and imbalance, revealing that a phase shift of the crystal structure is compensated by a superfluid counterflow. Such a phase shift provides evidence that the low-energy crystal Goldstone mode of the finite-size system is excited [14].
To probe other signatures of superfluidity, we investigate the collective breathing and dipole modes of the system, which we detect through the evolution of the cloud’s width $\sigma$ and the displacement of its center of mass $x_{\mathrm{0}}$, respectively. In conventional BECs, a ratio of their frequencies $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=\sqrt{5/2}$ (with $\omega_{\mathrm{B}}$ and $\omega_{\mathrm{D}}$ the breathing and dipole mode frequencies) signals a system described by superfluid hydrodynamics and thus supports the superfluid character of the cloud [34]. Fig. 3(b) shows the breathing oscillation of our system. We find a frequency of $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1$, in good agreement with the conventional BEC prediction. This behavior is consistent with numerical simulations of the Gross-Pitaevskii equation which implicitly assume superfluid hydrodynamics [33]. Here the dipole mode frequency $\omega_{\mathrm{D}}$ differs from the trapping frequency $\omega_{x}$ because spin-orbit-coupled systems violate Kohn’s theorem [35, 36], analogously to superfluids subjected to an optical lattice [37, 38]. Another interesting aspect is that, due to the changing size of the cloud, the number of stripes in the supersolid varies with time. The bottom panel of Fig. 3(b) shows that the superfluid flow can dynamically create additional stripes. Thus, our data support the superfluid character of the system via the frequency of its breathing mode and the existence of dynamic particle flow between the stripes.
IV.   Stripe compression excitations and location of the supersolid phase transition
Figure 4: Stripe compression mode and supersolid phase transition. (a) In situ observation of the stripe compression mode. After a rapid ramp of the coupling strength $\Omega$ ending at $t=0$, the spacing of the stripes $d$ oscillates in time, demonstrating that the crystal structure is not stiff. Errorbars denote one eom. Bottom: sketch of the stripe compression in real space (left) and in momentum space (right). (b) Observation of mode softening from momentum-space measurements. Top: collective frequencies of the dipole mode (red squares) and stripe compression mode (blue circles), in units of the trapping frequency $\omega_{x}$, vs. coupling strength $\Omega$. Solid lines: mixture model predictions without fitting parameters. Vertical black dashed line: thermodynamic limit prediction of the critical point $\Omega_{\mathrm{c}}$ [29]. Bottom: example oscillations of the relative momentum of the dressed states $\hbar\Delta k$ for $\hbar\Omega=0$, $1.50\pm 0.06$ and $2.10\pm 0.08\,E_{\mathrm{R}}$, with the exponentially decaying sinusoidal fits (solid lines) used to extract the corresponding frequencies. Errorbars: error on the fit (top panel) and standard deviation of up to three repetitions (bottom panel). (c) Top: frequency softening of the stripe compression mode for magnetic fields $B=51.52\pm 0.02$, $51.75\pm 0.02$ and $51.85\pm 0.02$ G (from dark to light blue, respectively). Bottom: phase diagram of the supersolid-to-plane-wave phase transition, calculated from the mixture model. Solid blue lines: magnetic field values corresponding to the top panel. Gray dashed line: phase boundary predicted by the mixture model.
Supersolids host a rich excitation spectrum originating from both their crystal-like and their superfluid character. Previous studies have revealed significant differences between platforms. Dipolar supersolids host crystal phonon-like modes [13, 15]. Instead, phonons are absent in cavity supersolids and multimode cavities are required to restore them [39]. This difference has led to controversial discussions on a supposed stiffness of the stripe pattern of spin-orbit-coupled supersolids, which were only recently clarified theoretically with the prediction of stripe compression modes akin to those of dipolar supersolids [26]. Here, we address this question experimentally. To this end, we prepare the system by a non-adiabatic ramp of the spin-orbit coupling strength $\Omega$, which is slow enough to remain in the lower dressed state, but sufficiently fast to excite the stripe pattern [29]. Figure 4(a) depicts the evolution of the modulation period $d$ over time for a system with $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$. Remarkably, we observe that $d$ is not constant, but instead displays a dynamical oscillation corresponding to a stripe compression mode. Its appearance proves that supersolids in spin-orbit-coupled BECs are not stiff. Observing this mode is a smoking gun of a fully dynamic supersolid, as it reveals the existence of a crystal phonon branch [26].
In a momentum-space picture, this mode is caused by an out-of-phase oscillation of the $\ell$ and $r$ condensates in their dispersion minima, and within our effective mixture model thus corresponds to a spin-dipole mode. This mode disappears at the transition to the plane-wave phase. Its frequency has been theoretically predicted to reduce its value when approaching the supersolid stripe to plane-wave phase transition, reaching a minimum at the transition point $\Omega_{\mathrm{c}}$ [40, 41]. In the next series of experiments, we therefore focus on determining the critical value $\Omega_{\mathrm{c}}$ for a spin-orbit coupling detuning $\delta=0$ through the frequency of the compression mode. To measure its softening, it is essential to prepare the ground state of the system. In spin-orbit-coupled BECs, this is challenging due to the existence of very long-lived metastable states [20] with spin polarization $P\not=P_{0}$, where $P_{0}$ is the polarization of the ground state. In the limit of large system sizes our mixture model predicts that, for each value of $\Omega$, $P_{0}$ minimizes the stripe compression mode frequency $\omega_{\mathrm{SC}}$ [29]. We thus experimentally minimize $\omega_{\mathrm{SC}}$ as a function of polarization $P$ for different coupling strengths $\Omega$, which allows us to investigate the expected mode softening of the compression mode when approaching $\Omega_{\mathrm{c}}$. To enhance our detection in the regimes of weak coupling $\Omega$ and large spin polarization $P$, we exploit momentum space images to perform this measurement [29].
Figure 4(b) summarizes our results. It displays the frequencies of the dipole and stripe compression modes as a function of the spin-orbit coupling strength $\Omega$, measured at the polarization $P$ that experimentally minimizes $\omega_{\mathrm{SC}}$, thus approximating the ground state polarization $P_{0}$. We observe that the frequency of both modes diminishes when increasing $\Omega$. For the dipole mode (red squares), it is due to the reduction of curvature of the dispersion relation at the minima (i.e., an increase of the effective mass [42, 43]). The frequency of the dipole mode vanishes at the plane-wave to single-minimum transition, which in the single-particle regime takes place at $\hbar\Omega=4\,E_{R}$ and has been investigated in previous works [35, 36]. It does not show particular features in the Raman coupling range explored here. In contrast, the frequency of the compression mode (blue circles), dramatically decreases when approaching the expected supersolid phase transition point $\Omega_{\mathrm{c}}$ (vertical black dashed line). The softening of the mode can also be clearly observed in the three exemplary oscillations included in the figure.
We compare our experimental results with our theoretical prediction (solid lines) and find a good agreement. The theoretical lines have no fitting parameters and result from a method originally developed to analytically predict the mode frequencies in conventional mixtures [44], which we adapt to our mixture model [29]. In the vicinity of the phase transition point the experimental data deviates from the thermodynamic-limit theory due to finite size effects, which give a finite gap to the zero-energy modes. All measurements displayed in Fig. 4(b) are taken at a magnetic field $B=51.75\pm 0.02$ G, resulting in a supersolid phase transition located at $\hbar\Omega_{\mathrm{c}}=2.05\pm 0.08\,E_{\mathrm{R}}$ (vertical dashed line). By exploiting the rich Feshbach spectrum of 41K [32], we can furthermore tune the bare-state interactions and displace the phase transition point. The top panel of Fig. 4(c) summarizes the results of such measurements, which are in good agreement with the theoretical expectations (solid lines) and the predicted phase diagram of the system as a function of the magnetic field (bottom panel). This final measurement showcases the tunability of the supersolid phase transition in our system.
V.   Conclusion
We have leveraged the tunable interaction properties of potassium atoms to realize a robust and high-contrast supersolid stripe phase in a spin-orbit-coupled BEC. We have imaged in situ its density modulation and have observed its rich dynamics. By observing a crystal compression mode, we have demonstrated that spin-orbit-coupled supersolid stripes are not stiff. We have used the frequency softening of this mode to locate the phase transition. Natural extensions of our work include the observation of other supersolid excitation modes, such as the crystal Higgs mode, and collective spin breathing [40, 33] and stripe angle oscillations [26], that could clarify how the spin degree of freedom further distinguishes spin-orbit-coupled supersolids from other platforms. Our system also opens the possibility of investigating the superfluid fraction across the phase transition [24]. Finally, by setting the effective interactions between the dressed condensates to attractive values ($a_{\ell r}<0$) it is possible to cancel the overall mean-field energy and reveal beyond mean-field effects. They are expected to stabilize quantum liquid droplets [45, 46] with supersolid properties [47, 48], thus providing a platform to systematically investigate the influence of quantum fluctuations on supersolidity.
Acknowledgments
We acknowledge insightful discussions with S. Stringari, K. T. Geier, G. I. Martone, P. Hauke, W. Ketterle, I. B. Spielman, F. Ferlaino, and T. Donner. We also wish to thank T. Donner, W. Ketterle, P. Massignan, A. Rubio-Abadal and S. Stringari for a critical reading of the manuscript.
Funding:
We acknowledge funding from the European Union (ERC CoG-101003295 SuperComp), the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (LIGAS projects PID2020-112687GB-C21 at ICFO and PID2020112687GB-C22 at UAB, and Severo Ochoa CEX2019000910-S), the EU QuantERA project DYNAMITE (funded by MICN/AEI/ 10.13039/501100011033 and by the European Union NextGenerationEU/PRTR PCI2022-132919 (Grant No. 101017733)), Deutsche Forschungsgemeinschaft (Research Unit FOR2414, Project No. 277974659), Generalitat de Catalunya (AGAUR SGR 2021-SGR-01448 at ICFO and 2021-SGR-00138 at UAB, CERCA program), Fundació Cellex and Fundació Mir-Puig. C.S.C., S.H., and R.R. acknowledge support from the European Union (Marie Skłodowska-Curie–713729, Marie Skłodowska Curie–101149245 Epiquant, and Marie Skłodowska-Curie–101030630 UltraComp). V.B.M. acknowledges support from the Beatriu de Pinós Program and the Ministry of Research and Universities of the Government of Catalonia (Grant No. 2019-BP-00228). J.C. acknowledges support from the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and A.C. acknowledges support from the UAB Talent Research program during the early stages of this work.
Author contributions:
C.S.C. and R.R. developed and took preliminary data on the first Raman coupling and matter-wave optics magnification experimental schemes. S.H., V.B.M. and R.V. developed the final magnification scheme, devised the experimental collective excitation methods and observables, and took the final data. S.H. and R.V. analyzed the final data. J.C., C.S.C, L.T., and A.C. developed the theory. J.C., C.S.C., and V.B.M. performed numerical simulations. S.H., R.V., J.C., and L.T. wrote the manuscript. L.T. developed the concept, and A.C. and L.T. supervised the work.

### v2

@@
 In particular, they host the so-called stripe phase, an instance of a supersolid state of matter.
-The peculiar excitation spectrum of the stripe phase, a definite footprint of its supersolidity, has so far remained out of experimental reach.
-We achieve in situ imaging of the stripes and directly observe both superfluid and crystal excitations.
+The peculiar excitation spectrum of the stripe phase, a definite footprint of its supersolidity, has been difficult to measure experimentally.
+Here, we perform in situ imaging of the stripes and directly observe both superfluid and crystal excitations.
 We investigate superfluid hydrodynamics and reveal a stripe compression mode, thus demonstrating that the system possesses a compressible crystalline structure.
 Through the frequency softening of this mode, we locate the supersolid transition point.
-Our results establish spin-orbit-coupled supersolids as a platform of choice to investigate supersolidity and its rich dynamics.
+Our results establish spin-orbit-coupled supersolids as ideal systems to investigate supersolidity and its rich dynamics.
 Supersolidity is an exotic phase of matter characterized by two spontaneously-broken continuous symmetries, translational symmetry and U(1) gauge symmetry [1, 2].
-They are associated to two independent order parameters and manifest themselves as a modulated density, reminiscent of a solid, and a global phase coherence, characteristic of a superfluid [3, 4, 5].
-This dual nature leads to a remarkably rich excitation spectrum with both superfluid and crystal excitations [6], including two gapless Goldstone modes.
-After the first realizations of supersolids in quantum gases [7, 8, 9, 10, 11], exploring their dynamics has thus been the focus of experimental research in recent years [12, 13, 14, 15, 16, 17, 18, 19] as it provides an unambiguous signature of supersolidity.
-Bose-Einstein condensates (BECs) with spin-orbit coupling, where the internal states of the particles are linked to their momentum through optical coupling [20], have been shown to host a supersolid state of matter in the form of the so-called stripe phase.
-Due to its conceptual simplicity, it can be described analytically and is thus ideally suited to study the exotic properties of supersolids [21, 22, 20, 23, 24].
-However, only very fragile realizations of the stripe phase have been demonstrated so far, and the contrast of the stripes, i.e., of the density modulation, was vanishingly small [8, 25].
-As a result, the excitations of this phase have been widely believed to be experimentally inaccessible and a debate has arisen over their nature [26].
-It was fueled by the observation that other platforms fall into two distinct categories: supersolids that support phonon modes [13, 15], and supersolids that are stiff [12], meaning that the periodicity of their density modulation is fixed.
-This led to the following question: are spin-orbit-coupled supersolid stripes stiff or do they host phonon modes?
-In this work, we experimentally explore the excitation spectrum of the stripe phase and show that spin-orbit-coupled supersolids have a compressible crystal structure and are not stiff.
-By harnessing the tunable interaction properties of ultracold potassium, we overcome the limitations of previous approaches and realize a stable stripe phase that persists at coupling regimes with strong modulation contrast.
-While the modulated density pattern could only be probed indirectly in previous experiments [8, 25], here we achieve in situ imaging of the stripes using a matter-wave magnification scheme.
-By exciting the collective modes of the system, we are able to address both superfluid and crystal-like aspects of supersolids.
-We show the existence of particle flow between the stripes and observe a stripe compression mode, i.e., an oscillation of the modulation period, which unambiguously proves the compressibility of the crystalline structure.
-Furthermore, we locate the phase transition from the supersolid to an unmodulated superfluid through frequency softening of the compression mode.
+They are associated with two independent order parameters and manifest themselves, respectively, as a modulated density, reminiscent of the crystalline structure of a solid, and a global phase coherence, characteristic of a superfluid [3, 4, 5].
+In recent years, several quantum-gas platforms have realized superfluids with a spontaneous density modulation [6, 7, 8, 9, 10].
+Among these, Bose-Einstein condensates (BECs) with spin-orbit coupling, where the internal states of the particles are linked to their momentum through optical coupling [11], have been shown to host a supersolid state of matter where the density modulation takes the form of stripes.
+Owing to its conceptual simplicity, this so-called stripe phase is analytically tractable and serves as a powerful platform for exploring supersolid phenomena [12, 13, 11, 14, 15].
+Figure 1: A robust high-contrast spin-orbit-coupled supersolid.
+(a) Spin-orbit-coupled BECs at small coupling have a dispersion relation with two minima, leading to two possible many-body phases: the unmodulated plane-wave phase (one minimum occupied) and the modulated supersolid-stripe phase (both minima occupied).
+(b) Two-photon Raman coupling scheme with two beams of frequency difference $\Delta\omega_{\mathrm{R}}$, coupling strength $\Omega$, detuning $\delta$, and recoil momentum per beam $\hbar k_{\mathrm{R}}$ (tunable through the angle of the Raman beams $\theta$), used to engineer spin-orbit coupling.
+(c) Dispersion relation $E(k)$ of the lower optically-dressed state (solid line), compared to that of the bare atomic states (dashed lines).
+Colorscale: spin polarization $P$ of the dressed state.
+The system can be modeled as a mixture of condensates in the left $\ell$ and right $r$ wells, which consist of a coherent superposition of the bare spin states (red and blue circles) and have modified effective interaction properties given by the effective scattering lengths $a_{\ell\ell}$, $a_{\ell r}$ and $a_{rr}$.
+(d) Effective interaction parameters at $\delta=0$ vs.
+Raman coupling $\Omega$, calculated via the mixture model.
+Vertical black dashed line: supersolid phase transition point $\Omega_{\mathrm{c}}$ separating the supersolid-stripe phase and the plane-wave phase.
+This panel shows how tuning interaction strengths can stabilize the stripe phase vs.
+the plane-wave phase.
+(e) Phase diagram of a homogeneous spin-orbit-coupled BEC featuring a large and stable supersolid stripe phase, calculated with the variational ansatz of [14, 16].
+$C$ is the contrast of the density modulation.
+Dashed lines: phase boundaries calculated with the mixture model.
+Gray band: region experimentally explored in this work, with detuning $\hbar\delta=0.00\pm 0.02E_{\mathrm{R}}$.
+Parameters in (d) and (e) correspond to 41K at a magnetic field $B=51.7$ G and with a density $1\text{\times}{10}^{14}\text{\,}\text{atoms}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$, which realizes a robust and high-contrast supersolid stripe phase.
+In supersolids, the existence of two order parameters leads to two gapless Goldstone modes and to a plethora of superfluid and crystal excitations [17].
+Probing this remarkably rich excitation spectrum has become a central topic of research across the different quantum-gas systems displaying supersolid properties [18, 19, 20, 21, 22, 23, 24, 25].
+Recently, the compressibility of the supersolids' emergent crystal structure has been put forward as an additional requirement for full-fledged supersolidity, given that solids naturally host crystal phonons.
+This property has been observed in the supersolid phase of dipolar quantum gases [19, 21, 23], whereas it is lacking in BECs where effective long-range interactions are induced by single-mode cavities [6, 18].
+Addressing this question in spin-orbit-coupled BECs has been challenging.
+Because of the extreme fragility of the realized stripe states, characterized by a vanishingly small stripe contrast [7, 26], the excitations of the supersolid phase in these systems have been considered experimentally inaccessible, and its properties have been widely debated [27].
+By utilizing spin-orbit-coupled potassium Bose-Einstein condensates, where the relative strengths of interspin and intraspin interactions are tunable through Feshbach resonances [28], we overcome the limitations of previous approaches and realize a stable stripe phase that persists at coupling regimes with strong modulation contrast.
+We are thus able to experimentally explore its excitation spectrum, show that its crystal structure is compressible, and precisely locate the supersolid phase transition.
 I. Supersolid spin-orbit-coupled condensates
-[…]
-In spin-orbit-coupled BECs, supersolidity stems from the unconventional dispersion relation and the momentum-dependent interactions that arise due to the mixing of spin and spatial degrees of freedom [21, 20, 22, 23].
-The dispersion relation displays two minima in momentum space [28], which results in two possible phases at the many-body level, sketched in Fig.
-1(a).
+In spin-orbit-coupled BECs, supersolidity stems from the unconventional dispersion relation and the momentum-dependent interactions that arise thanks to the mixing of spin and spatial degrees of freedom [12, 11, 13, 14].
+The dispersion relation displays two minima in momentum space [29], which results in two possible phases at the many-body level (Fig.
+1(a)).
 The plane-wave phase, which appears when a single minimum is occupied, has a homogeneous density profile.
-In contrast, the stripe phase, which occurs when both minima are occupied, has a spontaneously modulated density profile while it retains the spatial phase coherence of the condensate [21, 22, 20, 23], and constitutes a supersolid [6].
+In contrast, the stripe phase, which occurs when both minima are occupied, has a spontaneously modulated density profile while it retains the spatial phase coherence of the condensate [12, 13, 11, 14], which constitutes a supersolid [17].
 The experimental challenge is to create a situation where the stripe phase occupies a significant portion of the phase diagram [11].
@@
 This model considers the dressed states at the two minima of the dispersion relation, which emerges when two different states $\downarrow$ and $\uparrow$ (e.g., internal atomic states [11, 26], or different bands of an optical lattice [7]) are coupled through an optical field that transfers momentum to the atoms.
-A pair of laser beams in a two-photon Raman configuration with frequency difference $\Delta\omega_{\mathrm{R}}$, two-photon Rabi frequency $\Omega$ and two-photon detuning $\delta$ is used, as depicted in Fig.
-1(b).
-It leads to a momentum transfer $2\hbar k_{\mathrm{R}}$ along the $x$ direction $\mathbf{e_{x}}$, with $\hbar k_{\mathrm{R}}$ the recoil momentum unit and $\hbar=h/(2\pi)$ the reduced Planck’s constant.
+To realize this coupling, we use a pair of laser beams in a two-photon Raman configuration with frequency difference $\Delta\omega_{\mathrm{R}}$, two-photon Rabi frequency $\Omega$ and two-photon detuning $\delta$ (Fig.
+1(b)).
+It leads to a momentum transfer $2\hbar k_{\mathrm{R}}$ along the $x$ direction $\mathbf{e_{x}}$, with $\hbar k_{\mathrm{R}}$ the recoil momentum unit and $\hbar=h/(2\pi)$ the reduced Planck constant.
 The atom-photon dressed states have a momentum-dependent spin composition, which is set by the polarization parameter $P=\tilde{\delta}/\tilde{\Omega}$, where $\tilde{\delta}=\delta-2\hbar k_{\mathrm{R}}k/m$ and $\tilde{\Omega}=\sqrt{\Omega^{2}+\tilde{\delta}^{2}}$ are the generalized detuning and Rabi frequency.
@@
 $\mathcal{H}_{\mathrm{kin}}=\frac{\hbar^{2}}{2m}(\mathbf{k}+k_{\mathrm{R}}\sigma_{z}\mathbf{e_{x}})^{2}-\frac{\hbar\delta}{2}\sigma_{z}+\frac{\hbar\Omega}{2}\sigma_{x},$ (1)
-where $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices.
-For $\hbar\Omega<4\,E_{\mathrm{R}}$, where $E_{\mathrm{R}}=\hbar^{2}k_{\mathrm{R}}^{2}/2m$ is the Raman recoil energy, the dispersion relation of the lower dressed state features two minima, see Fig.
-1(c).
+where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli matrices.
+For $\hbar\Omega<4\,E_{\mathrm{R}}$, where $E_{\mathrm{R}}=\hbar^{2}k_{\mathrm{R}}^{2}/2m$ is the Raman recoil energy, the dispersion relation of the lower dressed state features two minima (Fig.
+1(c)).
 The system behaves as a mixture of condensates in the left ($\ell$) and right ($r$) wells that are non-orthogonal.
-Thus, a modulated density profile emerges due to interference between them when both $\ell$ and $r$ are simultaneously occupied and miscible [28, 20].
-Their overlap increases with the spin-orbit coupling strength, leading to a contrast of the density modulation that, away from the transition, scales linearly with $\Omega$ [29].
+Thus, a modulated density profile emerges owing to an interference between them when both $\ell$ and $r$ are simultaneously occupied and miscible [29, 11].
+As the spin-orbit coupling strength increases, $\ell$ and $r$ become less orthogonal to each other.
+This larger wavefunction overlap leads to an increased contrast of the density modulation which, away from the transition, scales linearly with $\Omega$ [30].
 Figure 2: In situ observation of supersolid stripes.
 (a) Single shot experimental image of the in situ density profile of the cloud at $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$ measured with matter-wave optics.
-The image is rescaled to represent the expected cloud aspect ratio before matter-wave magnification along $x$.
+The image is rescaled to represent the expected cloud aspect ratio before matter-wave optics magnification along $x$.
 Top right: corresponding density profile integrated along the stripes.
@@
 Raman coupling $\Omega$ measured from in situ (blue circles) and time-of-flight (gray squares, see [30]).
-Solid line: single-particle theory prediction, obtained from the inverse of the momentum difference of the minima of the dispersion relation (shown on the right for different spin-orbit coupling strengths $\Omega$).
+Solid line: single-particle theory prediction, obtained from the inverse of the momentum difference of the minima of the dispersion relation (shown in the inset for different spin-orbit coupling strengths $\Omega$) [30].
 Errorbars correspond to one standard error of the mean (eom).
@@
 Our effective field theory is based on the effective interactions of the optically-dressed states, which are given by the polarization at the dispersion minima [30].
-We name it the mixture model because it builds on the mixture picture introduced by [28, 20].
-It delivers analytic expressions for both ground state properties and collective excitations [29] and highlights the conceptual simplicity of spin-orbit-coupled supersolids.
-Figure 1(d) shows the effective interaction parameters according to our mixture model for a gas with bare scattering lengths $a_{\uparrow\uparrow}$, $a_{\downarrow\downarrow}$, and $a_{\uparrow\downarrow}$ at detuning $\delta=0$.
+We name it the mixture model because it builds on the mixture picture introduced in [29, 11].
+It delivers analytical expressions for both ground state properties and collective excitations [30] and highlights the conceptual simplicity of spin-orbit-coupled supersolids.
+Figure 1(d) shows the effective interaction parameters according to our mixture model for a gas with bare scattering lengths $_{\uparrow\uparrow}, aa_{\downarrow\downarrow}$, and $a_{\uparrow\downarrow}$ at detuning $\delta=0$.
 Increasing $\Omega$ reduces the effective intra-well interactions $a_{\ell\ell}$, $a_{rr}$ and increases the inter-well ones $a_{\ell r}$.
@@
 This situation is not available in atomic species like the 87Rb used in previous works, where the small scattering length difference yields a very fragile supersolid phase with a contrast of the density modulation $C<5\%$ [11, 31, 32, 26].
-However, we can achieve it by exploiting two-component potassium BECs, where the interactions are tunable using Feshbach resonances [32].
+However, we can achieve it by exploiting two-component potassium BECs, where the interspin interactions are tunable using a Feshbach resonance [28].
 The complete phase diagram of the system with 41K at 51.7 G in the parameter space of $\Omega$ and $\delta$ is depicted in Fig.
 1(e).
-The stripe phase appears as a dome that is not centered at $\delta=0$ due to the setting $a_{\downarrow\downarrow}<a_{\uparrow\uparrow}$.
+The stripe phase appears as a dome that is not centered at $\delta=0$ owing to the setting $a_{\downarrow\downarrow}<a_{\uparrow\uparrow}$.
 The phase diagram shows the contrast $C$ according to the variational model of Refs.
 [14, 16], in excellent agreement with the black dashed lines, which are analytical predictions from our mixture model.
-This constitutes an unusually large, stable, and high-contrast supersolid stripe phase and provides an excellent regime to perform our experiments.
+This constitutes an unusually large, stable, and high-contrast supersolid stripe phase and provides an excellent regime in which to perform our experiments.
 II. In situ observation of supersolid stripes
-We realize the supersolid stripe phase with a Bose-Einstein condensate of 41K atoms in states ||$\ket{\downarrow}\equiv\ket{F=1,m_{F}=0}$ and ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$.
-By setting the magnetic field to $B=51.7\pm 0.1$ G, between an intra-spin and an inter-spin Feshbach resonance, we tune the bare state interactions to $(a_{\uparrow\uparrow},a_{\downarrow\downarrow},a_{\uparrow\downarrow})=(115\pm 10,65\pm 1,40\pm 8)\,a_{0}$, where $a_{0}$ is the Bohr radius [32].
-We couple the two states with two Raman laser beams of orthogonal polarizations at the tune-out wavelength $\lambda_{\mathrm{R}}=768.97$ nm, for which scalar potentials cancel.
-We work at a zero two-photon Raman detuning $\hbar\delta=0.00\pm 0.02\,E_{\mathrm{R}}$ (gray band in Fig.
-1(e)), where the phase transition point is independent of the atomic density for our parameter regime [33].
-The Raman beams have a relative angle of $\theta=94\pm 1^{\circ}$, which leads to a momentum transfer along the $x$ direction of $2\hbar k_{\mathrm{R}}=4\pi\hbar\mathrm{sin}(\theta/2)/\lambda_{\mathrm{R}}$ [29].
+[…]
 […]
@@
 Figure 2(b) shows that the stripe spacing $d$ increases significantly with $\Omega$, which can be understood from a momentum-space picture.
-The stripe spacing is given by the momentum difference $\Delta k$ of the two dressed states through $d=2\pi/\Delta k$ which, at equilibrium, corresponds to the difference between the two minima of the dispersion relation (see right panel).
-It can also be inferred from spin-resolved momentum-space images obtained through time-of-flight expansion.
-Contrary to the in situ one, this analysis is model dependent [29], resulting in a spacing $\pi/k_{\mathrm{R}}$ at the zero-coupling point by construction.
-The data is shown as gray squares in Fig.
+The stripe spacing is given by the momentum difference $\Delta k$ of the two dressed states through $d=2\pi/\Delta k$ which, at equilibrium, corresponds to the difference between the two minima of the dispersion relation (Fig.
+2(b), inset).
+It can also be inferred from spin-resolved momentum-space images obtained through time-of-flight expansion (Fig.
+S4) [30].
+Contrary to the in situ analysis, the analysis of the time-of-flight data is model-dependent, resulting in a spacing $\pi/k_{\mathrm{R}}$ at the zero-coupling point by construction.
+The data are shown as gray squares in Fig.
 2(b).
@@
 Thus, our observations demonstrate that the period of spin-orbit-coupled supersolids is not externally imposed by the wavelength of the coupling laser.
-This variable crystalline pattern stands in contrast to cavity-mediated supersolidity, where the modulation period is fixed by the light wavelength [12].
+This variable crystalline pattern stands in contrast to that of cavity systems, where the modulation period is fixed by the light wavelength [18].
 III. Probing superfluid hydrodynamics through excitations
+Supersolids are characterized by a superfluid flow through the crystalline structure that maintains the global phase coherence of the system [5, 1].
+Whereas previous experiments demonstrated phase coherence in spin-orbit-coupled supersolids using Talbot interferometry [26], our in situ images allow us to investigate instead the superfluid hydrodynamic properties of the system.
+To this end, we followed the procedure developed for supersolids in dipolar gases [20] and studied the population imbalance between stripes.
+Figure 3(a) shows the imbalance between side stripes $\eta$ as a function of the displacement $D$ of the central stripe with respect to the center of mass of the cloud (see illustration of $D$ and $\eta$ in the inset and [30]).
+We see a clear linear correlation between displacement and imbalance, revealing that a phase shift of the crystal structure is compensated by a superfluid counterflow.
+Such a phase shift provides evidence that the low-energy crystal Goldstone mode of the finite-size system is excited [20].
 Figure 3: Crystal Goldstone mode and superfluid hydrodynamics.
@@
 displacement of the central stripe $D$ with respect to the center of mass of the cloud (see inset and [30]).
-For a harmonically-trapped system with superfluid flow, where a sliding of the density modulation is compensated by particle exchange between the stripes, both quantities are linearly correlated and provide evidence for the crystal Goldstone mode.
+For a harmonically-trapped system with superfluid flow, where a sliding of the density modulation is compensated by particle exchange between the stripes, the two quantities are linearly correlated and provide evidence for the crystal Goldstone mode.
 (b) Top panel: breathing mode of the overall cloud, observed through its size $\sigma$ vs.
@@
 Errorbars denote one standard eom.
-Supersolids are characterized by a superfluid flow through the crystalline structure that maintains the global phase coherence of the system [5, 1].
-While previous experiments demonstrated phase coherence in spin-orbit-coupled supersolids using Talbot interferometry [25], our in situ images allow us to investigate instead the superfluid hydrodynamic properties of the system.
-To this end, we followed the procedure developed for supersolids in dipolar gases [14] and studied the population imbalance between stripes.
-Figure 3(a) shows the imbalance between side stripes $\eta$ as a function of the displacement $D$ of the central stripe with respect to the center of mass of the cloud (see illustration of $D$ and $\eta$ in the inset and [29]).
-We see a clear linear correlation between displacement and imbalance, revealing that a phase shift of the crystal structure is compensated by a superfluid counterflow.
-Such a phase shift provides evidence that the low-energy crystal Goldstone mode of the finite-size system is excited [14].
-To probe other signatures of superfluidity, we investigate the collective breathing and dipole modes of the system, which we detect through the evolution of the cloud’s width $\sigma$ and the displacement of its center of mass $x_{\mathrm{0}}$, respectively.
+To probe other signatures of superfluidity, we investigate the collective breathing and dipole modes of the system, which we detect through the evolution of the cloud's width $\sigma$ and the displacement of its center of mass $x_{\mathrm{0}}$, respectively.
 In conventional BECs, a ratio of their frequencies $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=\sqrt{5/2}$ (with $\omega_{\mathrm{B}}$ and $\omega_{\mathrm{D}}$ the breathing and dipole mode frequencies) signals a system described by superfluid hydrodynamics and thus supports the superfluid character of the cloud [34].
-Fig.
-3(b) shows the breathing oscillation of our system.
+Figure 3(b) shows the breathing oscillation of our system.
 We find a frequency of $\omega_{\mathrm{B}}/\omega_{\mathrm{D}}=1.5\pm 0.1$, in good agreement with the conventional BEC prediction.
 This behavior is consistent with numerical simulations of the Gross-Pitaevskii equation which implicitly assume superfluid hydrodynamics [33].
-Here the dipole mode frequency $\omega_{\mathrm{D}}$ differs from the trapping frequency $\omega_{x}$ because spin-orbit-coupled systems violate Kohn’s theorem [35, 36], analogously to superfluids subjected to an optical lattice [37, 38].
-Another interesting aspect is that, due to the changing size of the cloud, the number of stripes in the supersolid varies with time.
+Here the dipole mode frequency $\omega_{\mathrm{D}}$ differs from the trapping frequency $\omega_{x}$ because spin-orbit-coupled systems violate Kohn's theorem [35, 36], analogously to superfluids subjected to an optical lattice [37, 38].
+Another notable aspect is that, owing to the changing size of the cloud, the number of stripes in the supersolid varies with time.
 The bottom panel of Fig.
@@
 Thus, our data support the superfluid character of the system via the frequency of its breathing mode and the existence of dynamic particle flow between the stripes.
-IV.   Stripe compression excitations and location of the supersolid phase transition
 Figure 4: Stripe compression mode and supersolid phase transition.
@@
 Gray dashed line: phase boundary predicted by the mixture model.
+IV. Stripe compression excitations and location of the supersolid phase transition
 Supersolids host a rich excitation spectrum originating from both their crystal-like and their superfluid character.
@@
 Dipolar supersolids host crystal phonon-like modes [19, 21].
-Instead, phonons are absent in cavity supersolids and multimode cavities are required to restore them [39].
-This difference has led to controversial discussions on a supposed stiffness of the stripe pattern of spin-orbit-coupled supersolids, which were only recently clarified theoretically with the prediction of stripe compression modes akin to those of dipolar supersolids [26].
+Instead, phonons are absent in cavity systems and multimode cavities are required to restore them [39].
+This difference has led to controversial discussions on the stiffness of the stripe pattern of spin-orbit-coupled systems, which were only recently clarified theoretically with the prediction of stripe compression modes akin to those of dipolar supersolids [27].
 Here, we address this question experimentally.
@@
 Figure 4(a) depicts the evolution of the modulation period $d$ over time for a system with $\hbar\Omega=2.00\pm 0.08\,E_{\mathrm{R}}$.
-Remarkably, we observe that $d$ is not constant, but instead displays a dynamical oscillation corresponding to a stripe compression mode.
-Its appearance proves that supersolids in spin-orbit-coupled BECs are not stiff.
-Observing this mode is a smoking gun of a fully dynamic supersolid, as it reveals the existence of a crystal phonon branch [26].
-In a momentum-space picture, this mode is caused by an out-of-phase oscillation of the $\ell$ and $r$ condensates in their dispersion minima, and within our effective mixture model thus corresponds to a spin-dipole mode.
-This mode disappears at the transition to the plane-wave phase.
-Its frequency has been theoretically predicted to reduce its value when approaching the supersolid stripe to plane-wave phase transition, reaching a minimum at the transition point $\Omega_{\mathrm{c}}$ [40, 41].
+Notably, we observe that $d$ is not constant, but instead displays a dynamical oscillation corresponding to a stripe compression mode.
+This proves that the stripe phase of a spin-orbit-coupled BEC has a compressible crystal structure.
+Observing this mode is a signature of full-fledged supersolidity, as it reveals the existence of a crystal phonon branch [27].
+In a momentum-space picture, this mode is caused by an out-of-phase oscillation of the $\ell$ and $r$ condensates in their dispersion minima.
+Within our effective mixture model, it corresponds to a spin-dipole mode of the dressed states.
+Owing to its spin nature, it decouples from collective modes that involve instead the density degree of freedom, such as the breathing mode observed in Fig.
+3(b) [30].
+This mode disappears at the transition to the plane-wave phase, although it can be observed in out-of-equilibrium configurations [40].
+Its frequency has been theoretically predicted to reduce when approaching the supersolid stripe to plane-wave phase transition, reaching a minimum at the transition point $\Omega_{\mathrm{c}}$ [41, 42].
 In the next series of experiments, we therefore focus on determining the critical value $\Omega_{\mathrm{c}}$ for a spin-orbit coupling detuning $\delta=0$ through the frequency of the compression mode.
 To measure its softening, it is essential to prepare the ground state of the system.
-In spin-orbit-coupled BECs, this is challenging due to the existence of very long-lived metastable states [20] with spin polarization $P\not=P_{0}$, where $P_{0}$ is the polarization of the ground state.
-In the limit of large system sizes our mixture model predicts that, for each value of $\Omega$, $P_{0}$ minimizes the stripe compression mode frequency $\omega_{\mathrm{SC}}$ [29].
-We thus experimentally minimize $\omega_{\mathrm{SC}}$ as a function of polarization $P$ for different coupling strengths $\Omega$, which allows us to investigate the expected mode softening of the compression mode when approaching $\Omega_{\mathrm{c}}$.
-To enhance our detection in the regimes of weak coupling $\Omega$ and large spin polarization $P$, we exploit momentum space images to perform this measurement [29].
+In spin-orbit-coupled BECs, this is challenging because of the existence of very long-lived metastable states [11] with spin polarization $P\not=P_{0}$, where $P_{0}$ is the polarization of the ground state.
+Such metastable states make it difficult to detect the phase transition point based only on the contrast of the stripes.
+Instead, in the limit of large system sizes our mixture model predicts that, for each value of $\Omega$, $P_{0}$ minimizes the stripe compression mode frequency $\omega_{\mathrm{SC}}$ [30].
+We thus experimentally minimize $\omega_{\mathrm{SC}}$ as a function of polarization $P$ for different coupling strengths $\Omega$, which allows us to investigate the expected mode softening of the compression mode when approaching the critical point $\Omega_{\mathrm{c}}$ (Fig.
+S5) [30].
+To enhance our detection in the regimes of weak coupling $\Omega$ or large spin polarization $P$, we exploit momentum space images to perform this measurement.
 Figure 4(b) summarizes our results.
@@
 We observe that the frequency of both modes diminishes when increasing $\Omega$.
-For the dipole mode (red squares), it is due to the reduction of curvature of the dispersion relation at the minima (i.e., an increase of the effective mass [42, 43]).
+For the dipole mode (red squares), this is caused by the reduction of curvature of the dispersion relation at the minima (i.e., an increase of the effective mass [43, 44]).
 The frequency of the dipole mode vanishes at the plane-wave to single-minimum transition, which in the single-particle regime takes place at $\hbar\Omega=4\,E_{R}$ and has been investigated in previous works [35, 36].
@@
 In contrast, the frequency of the compression mode (blue circles), dramatically decreases when approaching the expected supersolid phase transition point $\Omega_{\mathrm{c}}$ (vertical black dashed line).
-The softening of the mode can also be clearly observed in the three exemplary oscillations included in the figure.
+The softening of the mode can also be clearly observed in the three representative oscillations included in the figure.
 We compare our experimental results with our theoretical prediction (solid lines) and find a good agreement.
 The theoretical lines have no fitting parameters and result from a method originally developed to analytically predict the mode frequencies in conventional mixtures [45], which we adapt to our mixture model [30].
-In the vicinity of the phase transition point the experimental data deviates from the thermodynamic-limit theory due to finite size effects, which give a finite gap to the zero-energy modes.
+In the vicinity of the phase transition point the experimental data deviates from the thermodynamic-limit theory because of finite size effects, which give a finite gap to the zero-energy modes.
 All measurements displayed in Fig.
@@
 This final measurement showcases the tunability of the supersolid phase transition in our system.
-V.   Conclusion
-We have leveraged the tunable interaction properties of potassium atoms to realize a robust and high-contrast supersolid stripe phase in a spin-orbit-coupled BEC.
-We have imaged in situ its density modulation and have observed its rich dynamics.
-By observing a crystal compression mode, we have demonstrated that spin-orbit-coupled supersolid stripes are not stiff.
-We have used the frequency softening of this mode to locate the phase transition.
-Natural extensions of our work include the observation of other supersolid excitation modes, such as the crystal Higgs mode, and collective spin breathing [40, 33] and stripe angle oscillations [26], that could clarify how the spin degree of freedom further distinguishes spin-orbit-coupled supersolids from other platforms.
-Our system also opens the possibility of investigating the superfluid fraction across the phase transition [24].
-Finally, by setting the effective interactions between the dressed condensates to attractive values ($a_{\ell r}<0$) it is possible to cancel the overall mean-field energy and reveal beyond mean-field effects.
-They are expected to stabilize quantum liquid droplets [45, 46] with supersolid properties [47, 48], thus providing a platform to systematically investigate the influence of quantum fluctuations on supersolidity.
+V. Discussion and outlook
+Our realization of a high-contrast supersolid stripe phase in a spin–orbit-coupled BEC and the in situ observation of its crystal modes enables the observation of other excitations, in particular those involving the spin degree of freedom such as collective spin breathing [41, 33], stripe angle oscillations [27], and the Higgs amplitude mode of the translational symmetry breaking [18].
+Although the existence of a crystal Higgs mode is a fundamental property of supersolids, in dipolar systems it remains unobserved because it couples to higher excited modes [46, 47].
+In contrast, our mixture model suggests that in spin-orbit-coupled supersolids, the spin character of the Higgs mode keeps it decoupled from density excitations, making it more accessible.
+Besides their fundamental interest, investigating such excitations would clarify how the spin degree of freedom distinguishes spin-orbit-coupled supersolids from other platforms.
+Our system also opens the possibility of investigating the superfluid fraction across the supersolid phase transition [48], in analogy to very recent experiments with dipolar gases [24], but closer to the thermodynamic limit.
+Finally, by setting the effective interactions between the dressed condensates to attractive values ($a_{\ell r}<0$), it should be possible to cancel the overall mean-field energy and reveal beyond mean-field effects.
+These are expected to stabilize quantum liquid droplets [49, 50] with supersolid properties [51, 52], with lifetimes long enough to be observed.
+The ability to independently control the magnitude of beyond mean-field effects is a unique feature of spin-orbit-coupled supersolids, which should allow disentangling the effects of quantum fluctuations and supersolidity.
 Acknowledgments
@@
 Massignan, A.
-Rubio-Abadal and S.
+Rubio-Abadal, and S.
 Stringari for a critical reading of the manuscript.
 Funding:
-We acknowledge funding from the European Union (ERC CoG-101003295 SuperComp), the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (LIGAS projects PID2020-112687GB-C21 at ICFO and PID2020112687GB-C22 at UAB, and Severo Ochoa CEX2019000910-S), the EU QuantERA project DYNAMITE (funded by MICN/AEI/ 10.13039/501100011033 and by the European Union NextGenerationEU/PRTR PCI2022-132919 (Grant No.
+We acknowledge funding from the European Union (ERC CoG-101003295 SuperComp), the Spanish Ministry of Science and Innovation MCIU/AEI/10.13039/501100011033 (projects LIGAS PID2020-112687GB-C21, MAPS PID2023-149988NB-C22, and Severo Ochoa CEX2024-001490-S at ICFO and projects LIGAS PID2020-112687GB-C22 and MAPS PID2023-149988NB-C21 PID2020112687GB-C22 at UAB), the EU QuantERA project DYNAMITE (funded by MICN/AEI/ 10.13039/501100011033 and by the European Union NextGenerationEU/PRTR PCI2022-132919 (Grant No.
 101017733)), Deutsche Forschungsgemeinschaft (Research Unit FOR2414, Project No.
@@
 supervised the work.
+Competing interests:
+There are no competing interests to declare.
+Data and materials availability:
+The data shown in the main text and extended data figures are available from the CORA RDR repository [53].